OpenStudy (jordanloveangel):

Relationship B has a greater rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose all answers that are correct.

OpenStudy (jordanloveangel):

@Teddyiswatshecallsme @CausticSyndicalist @Kit_Kat_Nat_<3 @ameenbriscoe13 @nickmifsud1223 @Librarian @Mylove!

OpenStudy (jordanloveangel):

@perl @undeadknight26 @e.mccormick @dan815

OpenStudy (jordanloveangel):

CAN U GUYS HELP ME?

OpenStudy (anonymous):

ok

OpenStudy (jordanloveangel):

THX

OpenStudy (e.mccormick):

Well, do you see what the rate of change (slope) is on the graph of A?

OpenStudy (jordanloveangel):

idk how to slove it

OpenStudy (e.mccormick):

Do you know how to find a rate of change (slope) from a graph or from two points?

OpenStudy (jordanloveangel):

no

OpenStudy (e.mccormick):

Ever seen \(m=\dfrac{y_2-y_1}{x_2-x_1}\) aka: rise over run, aka chanhe in Y over change in X?

OpenStudy (jordanloveangel):

?

OpenStudy (e.mccormick):

How about \(y=mx+b\) ? All of these have slope in them. Usually you learn one or more of these before you get to relationships because a relationships is a specific type of function where as \(y=mx+b\) is just a generic line.

OpenStudy (jordanloveangel):

i get it lil

OpenStudy (e.mccormick):

OK. so, on any line, the slope is how fast it is going up or down: 

OpenStudy (e.mccormick):

In a relationship, the slope in the rate! So what they want is when the slope is higher than the slope of A. So what you need to find first is the slope of A.

OpenStudy (jordanloveangel):

ohh i see

OpenStudy (jordanloveangel):

these more then 1 right answer 4 this problem

OpenStudy (e.mccormick):

Possibly. Now, do you remember how to find slope from a graph? They gave you the graph of A for that reason.